# PlainQAFact: Retrieval-augmented Factual Consistency Evaluation Metric for Biomedical Plain Language Summarization

### v2

@@ Appendix A Dataset Annotation Protocol
Step 1: Sentence Annotation: Compared to the scientific abstract, analyze each sentence of the plain language summary across three dimensions: external information, category, and relation.
@@
Step 2: Using the Annotation Spreadsheet: You will work within a structured spreadsheet containing the segmented sentences from both the summaries and the scientific abstracts.
@@
<<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
@@
Corresponding Abstract <<s15_16>>For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low-certainty evidence). <<s15_17>>Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low-certainty evidence) and in time to defervescence (116 participants; one trial; low-certainty evidence). <<s15_20>>For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low-certainty evidence).
@@ Appendix B Dataset Examples
Appendix C LLM Prompts
@@
We utilize two types of prompts to guide LLMs through two stages of FC evaluation: classification and answer extraction. In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification. For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruc model.
@@
For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the FC evaluation performance of PlainQAFact. Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
@@
Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the FC of a given PLS based on its source scientific abstract.
@@
Example 4: Prompt of Llama 3.1 8B Instruct as a FC judge.
@@ 5. Negation
Appendix D Detailed Experiment Settings
All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory. We employ the Natural Language Toolkit (NLTK) [^5] https://www.nltk.org/ to split the plain language summaries into sentences using the punkt package. For the classifier fine-tuning, we set the random seed to 42 to split PlainFact into training, validation, and test sets. We tune the PubMedBERT-base model for three epochs with early stopping under the validation set. The final fine-tuning accuracy of the classifier on the test set is 0.77. The same seed (42) is also used to sample 200 summary-abstract pairs from the CELLS test set for the comparison in Table 2.
The QA model used in our PlainQAFact is downloaded from QAFactEval [39]. However, we set its maximum input length to 512 (from 364) tokens to incorporate more source context. Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens. In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (D).
We apply the default setting (i.e., RoBERTa-base) for AlignScore [19] and SummaC-Conv for SummaC [16]. For GPT-4o experiments as an factual consistency evaluator, we set the temperature as 1.0 in this paper.
Appendix E Statistical Testing
Following SummaC [16] and QAFactEval [39], we test whether our proposed metric achieves statistically significant improvements over other methods. Given Kendall’s $\tau$ and Pearson correlations do not show the discriminative feature of continuous scores in various evaluation metric, we run a systematic evaluation on four datasets (CELLS, FareBio, FactPICO, and PlainFact) through bootstrap resampling [56] on the AUC-ROC results. We compare our best metric to other methods using the confidence intervals as significance level of 0.05 and apply Bonferroni correction [57] following SummaC [16]. Our results show that PlainQAFact achieves statistically significant improvements over QuestEval [50], SummaC [16], and DAE [48] on the CELLS and PlainFact datasets. As shown in Figure 2, PlainQAFact also outperforms other automatic metrics in evaluating elaborative explanation sentences, except for QAFactEval [39]. This highlights the effectiveness of PlainQAFact for biomedical PLS tasks that include external explanations. On the other hand, when evaluating elaborative content in the FactPICO and FareBIO datasets (Figures 3 and 4), we do not find statistically significant differences across metrics. While some existing metrics perform similarly in certain settings, our metric consistently shows better performance across all three evaluation criteria.
Figure 3: Overall performance on summaries containing added information (i.e., elaborative explanations) from FactPICO [9]. The std. of PlainQAFact, Llama 3.1, and GPT-4o are 0.2, 0.2, and 3.3, respectively, based on five runs of each metric (details in  E).
Figure 4: Overall performance on summaries with elaborative explanations from FareBio [47]. The std. for PlainQAFact, Llama 3.1, and GPT-4o are 0.03, 0.3, and 5.7, respectively, computed over five runs (details in  E.)
Appendix F Summary-Level Explanation-Only Evaluation
For FareBio [47], we collect explanation sentences based on both “faithfulness” and “factual hallucination” labels in the original dataset. As shown in Figure 4, PlainQAFact achieves the best performance compared with all other metrics, showing a clear advantage over GPT-4o. Since the FareBio dataset provides sentence-level annotations of factual consistency for each sentence in the plain language summaries, the results are more reliable than those from FactPICO.
Appendix G Pilot Study on FactPICO

### v3

@@ Appendix A Dataset Annotation Protocol
Step 1: Sentence Annotation Instructions Compared to the scientific abstract, analyze each sentence of the plain language summary across three dimensions: external information, category, and relation.
@@
Step 2: Using the Annotation Spreadsheet You will work within a structured spreadsheet containing the segmented sentences from both the summaries and the scientific abstracts.
@@
Annotation Fields to Be Completed <<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
@@ Appendix B Dataset Examples
Corresponding Abstract <<s15_16>>For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low‐certainty evidence). <<s15_17>>Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low‐certainty evidence) and in time to defervescence (116 participants; one trial; low‐certainty evidence). <<s15_20>>For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low‐certainty evidence).
@@
Appendix C Qualitative Analysis of PlainFact
ID Target PLS Sentence A1 A2 Final Rationale
Ex1 Obesity is associated with many health problems and a higher risk of death. “Health problems” and “higher risk of death” are not in the original abstract. No association between obesity and health/death issues exists in the source.
Ex2 It can be hard for the person to gain control of the eye, which can cause distress in social situations. Most relevant source (from A1): “Spontaneously manifest DVD can be difficult to control and often causes psychosocial concerns.” However, no direct correlation between eye control and DVD exists; social situations are not equal to psychosocial concerns.
Ex3 Most participants were middle-aged. A2 provides supporting source: “Applicability of evidence was limited to middle-aged participants who were relatively free of co-morbidity and were undergoing elective abdominal surgery.” This sentence is supported by the source.
Ex4 In some people, winter blues becomes depression, which seriously affects their daily lives. A2 highlights relevant source: “SAD is a seasonal pattern of recurrent major depressive episodes that most commonly occurs during autumn or winter and remits in spring.” However, no causal relation between winter blues and depression in source.
Ex5 RDS is the most common cause of disease and death in babies born before 34 weeks’ gestation. Most relevant source: “RDS is the single most important cause of morbidity and mortality in preterm infants.” However, “babies born before 34 weeks’ gestation” is not mentioned in the abstract.
Table 8: Qualitative analysis of PlainFact annotation examples. $A1/A2$ = Annotator 1 or 2; ✓/✗ indicate yes or no judgments for the “Need External Information?” field.
Given the fair inter-rater agreement for the “Need External Information?” field in PlainFact, we conduct a targeted qualitative error analysis. We randomly sample one representative plain language sentence from each of five annotation batches where the two annotators disagree. Two annotators independently judge whether external information is required, and a third judge provides a brief rationale. As shown in Table 8, most disagreements fall into two categories: (i) differing interpretations of whether an elaborative explanation introduces new information beyond the source, and (ii) failures to locate or recognize supporting evidence in the source abstract. These examples highlight that the “Need External Information?” decision often depends on subtle judgments about sufficiency and specificity of source coverage, suggesting that clearer guidelines and additional training examples may further improve annotation consistency.
Appendix D Data and Indexing Statement
In this study, we follow the same protocol as used in MedRAG [37] to conduct sentence-level retrieval of the PLS sentences. Specifically, we use top-K as 3 across all of our experiments on both StatPearls [^9] https://huggingface.co/datasets/MedRAG/statpearls and medical Textbooks [^10] https://huggingface.co/datasets/MedRAG/textbooks [38] provided in the MedRAG repository [^11] https://github.com/Teddy-XiongGZ/MedRAG?tab=readme-ov-file#corpus. Specifically, the medical Textbooks corpus comprises 18 widely used medical textbooks (e.g., for USMLE preparation) originally collected by Jin et al. [38], while StatPearls corpus consists of point-of-care clinical decision support articles. The authors of MedRAG collect 9,330 publicly available articles from the NCBI Bookshelf [^12] https://www.ncbi.nlm.nih.gov/books/NBK430685/, which permits redistribution for non-commercial research purposes. The StatPearls corpus applies a hierarchical approach for chunking, treating each paragraph in an article as a snippet and splice all relevant hierarchical headings to form the corresponding title. The medical Textbooks is processed into chunks of no more than 1,000 characters. These corpus are made publicly available for research replication via the MedRAG HuggingFace repository [^13] https://huggingface.co/MedRAG under the original authors’ distribution terms. We do not redistribute raw StatPearls and Textbooks content beyond the MedRAG release and rely exclusively on their officially published corpus to ensure replicability. Researchers wishing to replicate the corpus index should (i) access the same StatPearls [^14] https://huggingface.co/datasets/MedRAG/statpearls and Textbooks [^15] https://huggingface.co/datasets/MedRAG/textbooks datasets, (ii) ensure their use complies with the copyright and permissions of the source license of MedRAG [37] and MedQA [38]. Following MedRAG, we use the same RecursiveCharacterTextSplitter from the LangChain [^16] https://www.langchain.com/ library to perform chunking. We directly retrieve top 3 most relevant snippets from both corpus. For the retriever, we use MedCPT [36] (specifically the ncbi/MedCPT-Query-Encoder [^17] https://huggingface.co/ncbi/MedCPT-Query-Encoder version, commit hash: d83a36c) for dense retrieval with CLS token pooling. Documents are indexed using FAISS [58] with an IndexFlatIP configuration (exact inner product search, , no quantization or clustering) for 768-dimensional embeddings. The retrieval system computes the inner product between query and document vectors to identify the most relevant snippets. This matches the default MedRAG implementation for MedCPT.
Appendix E Reproducibility
E.1 Fine-tuning QG Model
We fine-tune the QG model (i.e., BART-large) with the same datasets used in QAFactEval [39], including SQuAD [42] and QA2D [43]. We use a batch size of 16, a learning rate of 3e-5, maximum sequence length of 512 tokens, and two training epochs for QG model fine-tuning.
E.2 Fine-tuning Factuality Type Classifier
As mentioned in Section 3.2.1, we fine-tune a classifier using PubMedBERT-base model through our curated PlainFact dataset. Specifically, we set the random seed to 42 to split PlainFact into training, validation, and test sets. We tune the PubMedBERT-base model for three epochs with early stopping under the validation set. The final fine-tuning accuracy of the classifier on the test set is 0.77. The same seed (42) is also used to sample 200 summary-abstract pairs from the CELLS test set for the comparison in Table 2. The classifier used in PlainQAFact is fine-tuned on PlainFact with a batch size of 32 for 10 epochs. We apply early stopping during training based on validation loss.
E.3 General Details
For other open-source models used in this study, we apply the default version. For the QA model, we apply the original Electra-large model [^18] Can be downloaded here: https://github.com/salesforce/QAFactEval/blob/master/download_models.sh used in QAFactEval [39]. All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory. We employ the Natural Language Toolkit (NLTK) [^19] https://www.nltk.org/ to split the plain language summaries into sentences using the punkt package. For question filtering strategy, we follow the same protocol in QAFactEval. The temperature is set to 0 for GPT-4o and 0.01 for Llama 3.1. The Llama 3.1 model used in AE also uses a temperature of 0.01. The maximum input length is set to 512 for all models in PlainQAFact.The QA model (i.e., Electra-large) used in our PlainQAFact is directly downloaded from QAFactEval [39]. However, we set its maximum input length to 512 (from 364) tokens to incorporate more source context. Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens. In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (E). In the QF stage, we follow QAFactEval [39] and remove the unanswerable questions using the pre-trained Electra-large QA model [^20] Download here: https://github.com/salesforce/QAFactEval/blob/master/download_models.sh.
We apply the default setting (i.e., RoBERTa-base) for AlignScore [19] and SummaC-Conv for SummaC [16]. For GPT-4o experiments as a factual consistency evaluator, we set the temperature as 1.0 in this paper. For the FareBio dataset, each plain language sentence is annotated with two binary labels: faithfulness and factual hallucination. In the released data, there are no sentences that are faithful but factually wrong. For the dataset pre-processing, we treat a sentence as not factual only when both labels are “no.” If either faithfulness or factual hallucination is “yes,” we treat the sentence as factual. To construct summary level labels, we first apply this rule to every sentence and then group sentences by their original FareBio summary. If a summary contains at least one sentence that is not factual, we label the whole summary as not factual, which produces 33 summaries. If all sentences in a summary are factual, we label the summary as factual, which produces 174 summaries. For the subset of 53 valid summaries used in our explanation focused evaluation, we further keep only those sentences that meet our filtering criteria for factual elaborative explanations and then aggregate them back into summaries.
For the computational costs of GPT-4o, it costs about \$0.01 per plain language summary during the factual consistency evaluation on the CELLS dataset experiment, combining input and output tokens, and takes about 2 seconds per summary. In contrast, PlainQAFact takes about 30 seconds per summary on our hardware. However, PlainQAFact is able to provide intermediate outputs for each step in the QA based evaluation, which makes our metric more transparent and easier to inspect than the black box GPT-4o evaluation, even in most cases where GPT-4o achieves slightly higher performance than PlainQAFact.
E.4 LLM Prompts
@@
We utilize two types of prompts to guide LLMs through two stages of factual consistency evaluation: classification and answer extraction. In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification. For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruct model.
@@
For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the factual consistency evaluation performance of PlainQAFact. Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
@@
Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the factual consistency of a given PLS based on its source scientific abstract.
@@
Example 4: Prompt of Llama 3.1 8B Instruct as a factual consistency judge.
@@ 5. Negation
Appendix F Pilot Study on FactPICO
@@ F.2 Experiments and Results
Appendix G Statistical Testing
Following SummaC [16] and QAFactEval [39], we test whether our proposed metric achieves statistically significant improvements over other methods. Given Kendall’s $\tau$ and Pearson correlations do not show the discriminative feature of continuous scores in various evaluation metric, we run a systematic evaluation on four datasets (CELLS, FareBio, FactPICO, and PlainFact) through bootstrap resampling [59] on the AUC-ROC results. We compare our best metric to other methods using the confidence intervals as significance level of 0.01 and apply Bonferroni correction [60] following SummaC [16]. Our results show that PlainQAFact achieves statistically significant improvements over QuestEval [50], SummaC [16], and DAE [48] on the CELLS and PlainFact datasets. As shown in Figure 2, PlainQAFact also outperforms other automatic metrics in evaluating elaborative explanation sentences, except for QAFactEval [39]. This highlights the effectiveness of PlainQAFact for biomedical PLS tasks that include external explanations.
For hypothesis testing, we treat AUC-ROC as the primary metric. For each baseline $b$ we conduct a paired bootstrap test on the difference in AUC-ROC, $\Delta^{(k)}=\mathrm{AUC}^{(k)}_{\text{{PlainQAFact}}}-\mathrm{AUC}^{(k)}_{b}$ for bootstrap replicate $k$. The one-sided $p$-value for the alternative “PlainQAFact >> baseline” is estimated as $p=\Pr(\Delta\leq 0)$ over the 10,000 replicates. Across the seven baselines we apply Holm-Bonferroni [52] correction to control the family-wise error rate at $\alpha=0.01$.
Appendix H Entity Swap Perturbation
Dataset Metrics Kendall’s $\tau$ Pearson AUC (99% CI)
CELLS Llama 3.1 24.6 29.9 65.7* [57.8, 73.3]
GPT-4o 69.3 74.3 98.4 [96.4, 99.7]
QAFactEval 55.9 64.6 89.4 [85.5, 93.2]
QuestEval 22.6 27.0 66.0* [58.7, 72.8]
SummaC 16.4 24.5 61.7* [54.2, 68.7]
AlignScore 36.7 44.6 75.9* [69.5, 81.8]
DAE 20.6 20.6 56.8* [52.6, 60.9]
PlainQAFact 56.9 68.2 90.2 [85.9, 93.9]
Table 9: Evaluation results of automatic metrics on the CELLS dataset [28] with entity swap non-factual plain language summary perturbation. Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC (AUC-ROC). * indicates a statistically significant difference compared to PlainQAFact ($p<0.01$). The standard deviations (std.) over five runs are: 0.1 (PlainQAFact), 0.2 (Llama 3.1), and 2.5 (GPT-4o).
Dataset: CELLS AE Kendall’s $\tau$ Pearson AUC (99% CI) std.
w/ Heuristic Entity Swap Llama 3.1 56.9 68.2 90.2 [85.9, 93.9] 0.1
PromptRank 49.6 58.8 85.0* [79.7, 89.6] -
w/ LLM Perturbation Llama 3.1 62.1 75.2 93.8 [90.3, 96.7] 0.1
PromptRank 56.3 66.6 89.6* [85.3, 93.3] -
Table 10: Ablation study of answer extraction approach in the PlainQAFact on the CELLS dataset. w/ Heuristic Entity Swap uses APPLS entity swap [12] approach to create non-factual plain language summaries, while w/ LLM Perturbation applies GPT-4o to perturb the plain language summaries. AE = answer extraction. AUC = AUC-ROC. * indicates a statistically significant difference compared to Llama 3.1 ($p<0.01$). We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
As we discussed in Section 5.1, in addition to evaluating the performance of automatic factual consistency evaluation metrics on LLM-perturbed datasets, we also use the non-LLM entity swap approach introduced in Guo et al. [12] to assess the robustness of our proposed metric. Specifically, we perturb plain language summaries using the CELLS dataset following the same entity swap protocol as APPLS [12]. The entity swap system identifies specific “entity spans” in the text and substitutes them with existing terms from the UMLS knowledge base. As shown in Table 9, PlainQAFact achieves the best performance among all automatic evaluation metrics on the CELLS dataset. PlainQAFact achieves the highest correlation with human judgement (AUC=90.2, 99% CI [85.9, 93.9]). As comparative effect sizes, we report pairwise differences between PlainQAFact and the strongest baseline QAFactEval. The $\Delta\mathrm{AUC}=0.7$, where $\Delta$ is PlainQAFact minus the baseline. After correction, PlainQAFact shows significantly higher AUC than Llama 3.1, QuestEval, SummaC, AlignScore, and DAE ($p<0.01$, Holm-Bonferroni corrected for all), but is not significantly different from QAFactEval ($p=0.64$) or GPT-4o ($p=1.0$).
We also report an ablation study using a non-LLM answer extractor to show the effectiveness of our LLM-powered answer extraction approach. In Table 10, our answer extraction method using Llama 3.1 consistently achieves higher correlation (AUC=90.2, 99% CI [85.9, 93.9] in w/ heuristic entity swap and AUC=93.8, 99% CI [90.3, 96.7] in w/ LLM perturbation) compared with the non-LLM answer extractor (i.e., PromptRank). Using Llama 3.1 shows significantly higher AUC than PromptRank in both settings ($p<0.01$, Holm-Bonferroni corrected for all), highlighting the robustness and effectiveness of LLM-powered answer extraction in factual consistency evaluation.
Appendix I Summary-Level Explanation-Only Evaluation
Figure 5: Overall performance on summaries containing added information (i.e., elaborative explanations) from FactPICO [9]. The std. of PlainQAFact, Llama 3.1, and GPT-4o are 0.2, 0.2, and 3.3, respectively, based on five runs of each metric.
Figure 6: Overall performance on summaries with elaborative explanations from FareBio [47]. The std. for PlainQAFact, Llama 3.1, and GPT-4o are 0.03, 0.3, and 5.7, respectively, computed over five runs.
For FareBio [47], we collect explanation sentences based on both “faithfulness” and “factual hallucination” labels in the original dataset. As shown in Figure 6, PlainQAFact achieves the best performance compared with all other metrics, showing a clear advantage over GPT-4o. Since the FareBio dataset provides sentence-level annotations of factual consistency for each sentence in the plain language summaries, the results are more reliable than those from FactPICO. On both datasets, we do not find statistically significant differences in AUC-ROC metric by $p<0.01$. While some existing metrics perform similarly in certain settings, our metric consistently shows better performance in correlation.
Appendix J Quantitative Error Analysis of FareBio
Given the Table 6, we further conduct quantitative analysis regarding failure cases in FareBio using our proposed metric. We randomly sample 30 plain language summaries (both factual and non-factual) from FareBio and evaluate all sentences within them, resulting in 324 split sentences. We categorize errors into three main types:
1. QA Pipeline Errors (34.0% of errors): misclassification of sentence types, retrieval of irrelevant context, and generation of unanswerable questions.
2. Answer Extraction and Evaluation Issues (18.4% of errors): failure to extract numerical values and answer mismatches when using BERTScore.
3. Dataset Labeling Issues (47.7% of errors): inconsistencies in FaithBio’s annotation protocol.
Table 11 presents the complete quantitative breakdown. We identify 109 errors across the 324 sentences (33.6% error rate). Notably, 47.7% of errors occurs in sentences mislabeled as factual and non-factual, with most being false negatives. FareBio uses a different annotation protocol than PlainFact. All source-grounded simplification sentences are implicitly treated as factual. Annotators do not evaluate factuality once a sentence is labeled as faithful. This creates label noise when evaluating simplification sentences, as the ground truth may not reflect actual factual consistency. Our analysis reveals that PlainQAFact performs better on PlainFact (where annotation protocols align with our metric’s design) than on FareBio. The performance gap reflects both limitations in our approach, particularly for numerical verification, and incompatibilities between FareBio’s annotation protocol and our QA components.
Error Type Count % of Errors % of Total In F In NF
QA Pipeline Errors
Misclassification 13 11.9% 4.0% 10 3
Retrieval quality 9 8.3% 2.8% 8 1
Unanswerable generated question 15 13.8% 4.6% 11 4
Answer Extraction and Evaluation Issues
Numerical value mis-extraction 16 14.7% 4.9% 11 5
Answer match misalignment 4 3.7% 1.2% 3 1
Dataset Issues
False positive (F labeled as NF) 10 9.2% 3.1% 0 10
False negative (NF labeled as F) 42 38.5% 13.0% 42 0
Total Errors 109 100.0% 33.6% 85 24
Correct Classifications 215 66.4% 114 101
Total Sentences Analyzed 324 100.0% 198 126
Table 11: Quantitative error analysis of PlainQAFact on 60 randomly sampled summaries from FareBio dataset (30 factual, 30 non-factual). Summaries are split into sentences when evaluating the factual consistency. F = factual sentences; NF = non-factual sentences.